1. You are working in the Emergency Room and haloperidol 7.5 mg IM is ordered in agitated patients with schizophrenia.

Label reads;

Haloperidol 50mg/10mL

How many mL will you give?_______mL (answer to the nearest tenth of an mL)

2. The physician has orders for a patient with generalized anxiety disorder lorazepam (Ativan) 0.02 mg/kg every 3 hours IV PRN. The patient weight is 55kg. Lorazepam is available as 2mg/mL. How many mg do you give for each dose? ____mg. (Round to the nearest 10th)
Answer #1

ANSWER.

* Calculation of ml of drug required.

        Formula

              Volume(ml) = Desired dose/Dose in hand * Quantity.

   Here

       Desired dose (D) = 7.5 mg.

       Dose in hand (H) = 50 mg.

           Quantity (Q) = 10 ml

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 7.5/50 * 10

                                                        = 0.15 * 10

                                                        = 1.5 ml.

2) Calculation of mg give for each dose.

                   Ordered dose = 0.02 mg/kg.

              Weight of the patient = 55 kg.

     Amount(mg) of drug give per dose = 0.02 * 55

                                                               = 1.1

                                                             = 1 mg.